OpenStudy (anonymous):

what is the slope of the line given by the following equation? y=-x+4

OpenStudy (anonymous):

-1

OpenStudy (saifoo.khan):

y = mx + b Where "m" is the slope.

Directrix (directrix):

-1 because y = (-1) x + 4 is in the y = mx+b form. m, slope, = -1

OpenStudy (anonymous):

this can be solved in two ways input two values into x and record the y values that come out to acquire two points. Then use the equation slope = (y2 - y1)/(x2 - x1) Where (x1,y1) is the point furthest to the left on the x axis The second way this can be solved is using y=mx+b y = mx + b m = slope b = y-intercept x = input y = output find the y-intercept of the equation by setting x = 0 input the y value you obtain into b of the equation as it is your y-intercept plug an x value into the equation and collect a y value so you acquire a point insert that point into your y=mx+b equation and solve for m